On January 1, 2020, Windsor Company purchased the following two machines for use in its production process. Machine A: The caCalculate the amount of depreciation expense that Windsor should record for Machine B each year of its useful life under theWhich method used to calculate depreciation on Machine B reports the highest amount of depreciation expense in year 1 (2020)?
Answer #1
Machine A:
Total Cost of Machine A:
purchase cost 43500
sales tax 3550
shipping 150
Insurance during shipping 50
Installation and testing 200
Total Cost of Machine A 47450
Depreciation per year on straight line = (Cost - salvage)/life years
(47450 - 5250)/5 = 8440
Journal entry:
Date Accounts Title Debit $ Credit $
1-Jan-20 Machine A 47450
Cash 47450
(total cost of machine booked)
31-Dec-20 Depreciation exp. 8440
Acc. Dep. - Machine A 8440
(being 2020 depreciation booked)
Machine B:
Method : Straight line, Annual depreciation rate = 25%
Annual depreciation = (180000 - 10000)/4 = 42500
Year Depreciation exp. Acc. Dep. Book Value
1-Jan-20 180000
31-Dec-20 42500 42500 137500
2021 42500 85000 95000
2022 42500 127500 52500
2023 42500 170000 10000
Method: Declining balance - double, Dep. Rate = 25%*2 = 50%
Year Depreciation exp. Acc. Dep. Book Value
1-Jan-20 180000
31-Dec-20 90000 90000 90000
2021 45000 135000 45000
2022 22500 157500 22500
2023 11250 168750 11250
Method: Unis of activity method
Depreciation per unit = (180000-10000)/102000 = 1.666667
Year Units Depreciation exp. Salvage
2020 43000 71667
2021 30000 50000
2022 17000 28333
2023 12000 20000 10000
102000 170000
Depreciation expense:
Method 2020 2021 2022 2023 Total
Straight line 42500 42500 42500 42500 170000
Declining balance 90000 45000 22500 11250 168750
Units of activity 71667 50000 28333 20000 170000
Highest depreciation amount 2020
90000 (under Declining balance )
Highest amount of depreciation in year 4 (2023):
42500 (straight line)
Highest total amount over the 4-year period:
170000 (under straight line and Units of activity)